What are the x- and y-intercepts of the ellipse x^2/25+y^2/4=1
x = ±2, y = ±5
x = ±4, y = ±25
x = ±5, y = ±2
x = ±25, y = ±4



Answer by Alan3354(69443)

What are the x- and y-intercepts of the ellipse x^2/25+y^2/4=1
------------------
The x-intercepts are at y = 0:
x^2/25 = 1
---------------
The y-intercepts are at x = 0:
y^2/4 = 1